OpenStudy (mwhit2101):

SIXTH GRADE MATH HELP! MEDAL + FAN The graph of which function passes through the point (3, 21)?

OpenStudy (mwhit2101):

A. y=2x^2+3 B. y=-2x^2+3 C. Y=2x^2-3 D. Y=3x^2+2